# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v3

@@ Abstract.
With more than 11 times as many pageviews as the next largest edition, English Wikipedia dominates global knowledge access relative to other language editions. Readers are prone to assuming English Wikipedia as a superset of all language editions, leading many to prefer it even when their primary language is not English. Other language editions, however, comprise complementary facts rooted in their respective cultures and media environments, which are marginalized in English Wikipedia. While Wikipedia’s user interface enables switching between language editions through its Interlanguage Link (ILL) system, it does not reveal to readers that other language editions contain valuable, complementary information. We present WikiGap, a system that surfaces complementary facts sourced from other Wikipedias within the English Wikipedia interface. Specifically, by combining a recent multilingual information-gap discovery method (Samir et al., 2024) with a user-centered design, WikiGap enables access to complementary information from French, Russian, and Chinese Wikipedia. In a mixed-methods study (n=21), WikiGap significantly improved fact-finding accuracy, reduced task time, and received a 32-point higher usability score relative to Wikipedia’s current ILL-based navigation system. Participants reported increased awareness of the availability of complementary information in non-English editions and reconsidered the completeness of English Wikipedia. WikiGap thus paves the way for improved epistemic equity across language editions.
@@ 1. Introduction
Wikipedia boasts more than 300 language editions, inviting contributions from all over the world. The investment in creating multiple language editions reflects Wikipedia’s principles of democratizing both access to and contribution of information. By design, Wikipedia editions are not translations of one another, but rather reflect the specific knowledge that is relevant for their respective readers, often based on cultural experience and local media coverage (Avieson and Mazumdar, 2025; Kumar, 2021). As a result, readers of different Wikipedia editions are not exposed to the same information.
English Wikipedia in particular is the largest edition, which is read and edited by more people than any other language edition. This is perhaps unsurprising given that English is the major global lingua franca. However, given the complementary nature of Wikipedia editions, even readers of English Wikipedia don’t have access to content that may yet be relevant from other language editions.
One might argue that since English is a global language, editors from other linguistic regions can add culturally-relevant information into English Wikipedia. Indeed, prospective editors, knowing that English Wikipedia attracts a higher readership than any other edition, often opt to try to contribute to English Wikipedia, even when they possess a high degree of literacy in other languages (Hale, 2014). In 2024, English Wikipedia received nearly half of the 109M contributed edits (Wikipedia contributors, 2025).
In practice, however, the overwhelming dominance of English poses a significant challenge to egalitarian knowledge curation. It is not uncommon for such editors to contribute culture- or region-specific knowledge to English Wikipedia to encounter hegemonic resistance. While Wikipedia is in principle open to contributions from any person with literacy skills and an internet connection, it was reported in (2017) that an overwhelming majority of edits are contributed by a small portion of the editor base (Oberhaus, 2017). As of (2024b), for English Wikipedia, these editors were overwhelmingly male and white (Wikimedia Foundation, 2024b). This has resulted in contributions being rejected, even when the topic is geographically positioned in a non-Anglophone country; see, for example, a case study on a lengthy editor dispute resulting in the discarding of Indian editors’ cultural knowledge of the Ganga river in India (Kumar, 2017).
Yet, at the same time, English Wikipedia is by far the most widely read and authoritative source among the $300+$ language editions. In a 2011 report, $93\%$ of all editors surveyed attested to reading English Wikipedia, even though it was a primary language for only $52\%$ of them (Kumar, 2021; Wikimedia Foundation, 2011). We are thus presented with a remarkable asymmetry – English Wikipedia marginalizes potential contributions from non-Anglophone sources, yet inadvertently monopolizes attention from information-seeking readers away from language editions that contain these marginalized perspectives (e.g., Avieson and Mazumdar, 2025). This epistemic assumption of the completeness of English Wikipedia, which gives rise to this asymmetrically large readership, was previously aptly named the English-as-superset hypothesis (Hecht, 2013).
A promising approach for mitigating the hegemonic dominance of English Wikipedia is to increase engagement with information in other language editions. While these are smaller, they are by no means lacking in thoughtfully curated and often complementary content. As Hecht and Gergle (2010) showed, the English language edition was lacking coverage entirely for a considerable number of topics, with English Wikipedia lacking $60\%$ of the articles present in the Ukrainian edition, for instance. More recently, Samir et al. (2024) showed large information gaps in English Wikipedia compared to Russian and French Wikipedia across more than $2.5k$ Wikipedia biography pages, finding that English Wikipedia can lack even $50-70\%$ of content in other language editions when looking at regionally prominent public figures. Prior studies have sought to capitalize on these complementary multilingual texts, either through consolidation into a monolithic knowledge base or through interface redesigns. Both of these approaches, which we lay out in greater detail in Section 2, face fundamental obstacles in smoothly integrating multilingual content into readers’ engagement with English Wikipedia.
@@
In this work, we present WikiGap, a socio-technical system that surfaces sentence-level factual differences across language editions of Wikipedia. The system combines a computational backend (InfoGap) that identifies multilingual content gaps (Samir et al., 2024) with a browser-based interface that presents these differences through in-page highlights and a sidebar display (Figure 1). We address two research questions: First, how can multilingual content be seamlessly integrated into the English Wikipedia reading experience? Second, how might readers benefit from engaging with this cross-lingual content? To explore these questions, we developed a web extension that unobtrusively augments English Wikipedia articles with supporting facts drawn from the respective articles in the French, Russian, and Chinese editions.
@@
We evaluated WikiGap through a mixed-methods user study using a fact-finding task focused on culturally rich food articles – a domain where cross-lingual disparities are particularly salient (Laufer et al., 2015). Our results show that participants strongly preferred WikiGap over Wikipedia’s default interlanguage links system (ILL), with System Usability Scale (SUS) scores 32 points higher on average. In terms of performance, participants were significantly more accurate and faster at answering factual questions when using WikiGap, suggesting that surfacing multilingual information directly within the English interface not only improves user satisfaction but also enhances learning efficiency.
@@
Through a Post-Study interview, we find that participants, prior to using WikiGap, were largely unaware of content gaps in English Wikipedia, demonstrating that the problematic English-as-superset epistemic assumption persists more than a decade after it was identified in seminal studies by Hecht (2013). Participants were overwhelmingly appreciative of WikiGap’s capacity to surface these substantive information gaps during the reading of English Wikipedia pages. Moreover, a few participants with prior editing experience described how the system could support their editorial workflows by enabling source verification and facilitating the integration of multilingual content. These findings underscore WikiGap’s broader value as a bridge between communities of editors and readers across languages. At the same time, a minority of responses reflected concerns about the emergence of large language models (LLMs) in drawing readers away from information-seeking directly on Wikipedia, raising important questions about the increasingly marginalized role of provenance in accessing information, which should be investigated in future research. Overall, participants largely expressed openness and enthusiasm about interacting with multilingual content in English Wikipedia. We hope that this work paves the way for further growth of language editions that have been marginalized relative to English since Wikipedia’s inception in 2001 (Kumar, 2021). [^1] We are sharing the code and data: https://github.com/aw814/WikiGap
Our contributions are as follows:
  • A novel system design that leverages a computational pipeline for detecting fact-level information gaps across languages, integrating facts sourced from other language editions into English Wikipedia articles.
  • An empirical evaluation using a mixed-methods study (N=21) showing that WikiGap significantly improves fact-finding accuracy and completion time, achieves higher usability than Wikipedia’s default ILL system, and fosters greater engagement with multilingual content, thereby challenging the perceived epistemic supremacy of English Wikipedia.
@@ 2. Background and Related Work
Wikipedia is a peer production system (Benkler et al., 2015), where a large number of collaborators come together to produce a shared outcome. There are important norms in this collaborative practice that are referenced heavily in deliberating content contributions, although the explication and invocation of these norms can vary widely across language editions (Hwang and Shaw, 2022; Bipat et al., 2021). Community-developed tools such as vandalism patrol and bot-assisted help sustain these collaborative norms (Geiger and Ribes, 2010). Wikipedia has also introduced formal mechanisms like page protection policies, which increase friction in contributing edits to popular or controversial pages (Ajmani et al., 2023b). But the overarching ethos is that there should be no hard-set rules, and Wikipedia should remain largely self-organizing. These tensions between coordination, conflict, and self-organization mirror challenges seen across peer-production communities beyond Wikipedia, in thousands of other online wiki communities, for instance Wookiepedia and the Psychology Wiki (Kittur and Kraut, 2010). Large-scale collaborative peer production systems like Wikipedia are intended to democratize access to information, making it “easily available to people who would have otherwise had to seek it out in piecemeal and often haphazard fashion, if they could access it at all” (Justice and Golumbia, 2024, pp. 273). It is not simply about access either; Wikipedia is largely associated with an air of democratizing the production of knowledge, a telos of rebelling against centralized authority and bureaucracy. This more laissez-faire form of collaborative content production is underpinned by the idea that a distributed form of knowledge curation could result in egalitarian outcomes in the access to and production of knowledge.
But this vision has been argued to be highly idealistic, as the capacity to create archival encyclopedia content has long been recognized as sociopolitically contingent (Foucault, 2013; Said, 1977). For example, the composition of the English Wikipedia editor base (white, male) has predictably led to the marginalization of women on Wikipedia (Hargittai and Shaw, 2015; Wagner et al., 2016), as well as scholars from racial minority backgrounds (Adams et al., 2019). The demographic skews in the editor base are not merely incidental, but they can systematically shape whose knowledge is valued in the encyclopedia, privileging some forms of knowledge while marginalizing others. This can lead to reproducing epistemic injustice (Ajmani et al., 2024).
When it comes to multilingualism in Wikipedia, Wikipedia’s pillar for verifiable information, a component of the Neutral Point of View policy, has introduced barriers for creating content in other language editions. English is widely recognized as the preeminent vehicle for conveying scholarship, making it challenging to find primary sources for other language editions unless editors are fluent in English (Sen et al., 2015). While there are multilingual editors who contribute to multiple language editions, who serve as a bridge connecting different language editions (Hickman et al., 2021), they are a smaller proportion of editors on Wikipedia (Hale, 2014).
In this section, we highlight the perceived hegemonic status of the English language edition, while maintaining that it has information gaps that illustrate that it should not serve as the authoritative language edition (Sec. 2.1). We then discuss previous solutions, including a unified language edition, Wikipedia’s Interlanguage Link navigation panel, and other interfaces proposed for interacting with content in different language editions (Sec. 2.2). We argue that these prior works have yet to achieve an important goal of increasing readers’ interaction with content from other language editions. In achieving this goal, we challenge the perceived completeness of the English language edition—what Hecht (2013) has previously called the English-as-superset hypothesis—an important step towards greater democratization of knowledge production. In Sec. 3, we introduce our system, WikiGap, that augments the experience of reading English Wikipedia documents by showcasing information gaps relative to other language editions.
2.1. English Wikipedia is Considered the Default, yet Information Gaps with Other Editions Persist
In 2008, Wikimedia Foundation cofounder Jimmy Wales wrote that Wikipedia was for creating a free encyclopedia and making it available to “every single person on the planet in their own language” (Cohen, 2008). Unsurprisingly, considering the status of English as the global lingua franca, the English edition of Wikipedia has significantly outpaced other language editions in terms of size and coverage.
English Wikipedia is considered the “default” edition.
It is often taken for granted that English Wikipedia serves as a superset of the knowledge encoded in other language editions. Hecht (2013) identified that this “English-as-superset” viewpoint. Although not explicitly stated as such, this viewpoint is nonetheless remarkably prevalent, even among researchers who rely on multilingual Wikipedia to train text-generation and text-embedding models.
Knowledge gaps in English Wikipedia.
Hecht and Gergle’s ((2010)) formative study, however, found English Wikipedia was far more of a complement rather than a superset. They found considerable gaps in topic coverage in English Wikipedia relative to other language editions. Samir et al. ((2024)) further showed that even when a topic is covered across languages, English Wikipedia often lacks a large body of facts present in other editions, studying French and Russian Wikipedia specifically.
Knowledge is often culturally-positioned.
One might interpret Wales’ statement to mean that there exists a set of facts constituting a language-agnostic encyclopedia, that can then be cryptographically encoded into various languages. More generally, this view of language as an object that can exist outside of socio-geographic context is highly prevalent in philosophies of language, where language is metaphorized as a cryptographic encoding and the language-processing brain as (often literally) a computer (Golumbia, 2009, Chapter 4).
While some facts are culturally invariant, such as physical laws, some of Wikipedia’s most-read articles pertain to socio-geographically positioned histories of peoples, places, recipes, and so on (Wikimedia Foundation, 2024a). Generally speaking, all of Wikipedia’s language editions are necessarily socio-geographically positioned (Johnson and Lescak, 2022) and exhibit a “self-focus bias(Hecht and Gergle, 2009), the tendency of editors to contribute more to articles relevant to their region, including articles on people (Samir et al., 2024; Callahan and Herring, 2011), places (Oeberst and Ridderbecks, 2024), and cuisines (Laufer et al., 2015).
No language edition is complete.
For those culturally-positioned topics, the presupposition of a universal language-agnostic knowledge base has contributed to the marginalization of other language editions. English Wikipedia is treated as the default, standard, or “invisible center(Benjamin, 2023) through which other language editions are compared against. For example, Hickman et al. (2021) documents that bilingual Indian and Pakistani editors refer to English Wikipedia to determine the contributions to make to their local language editions, perceiving English Wikipedia to be the objective standard. [^2] This isn’t exclusive to South Asian editors; as we show in Section 5, readers and editors alike remain largely unaware of cross-linguistic heterogeneity in Wikipedia. In reality, neither the English version nor the local language edition for a topic is a complete article; a considerable proportion of unique facts can still be identified in non-local language editions (Samir et al., 2024). There is thus an incentive to look at multiple language editions, even when a topic has a strong regional association with one language edition.

### v4

@@ Abstract.
With more than 11 times as many pageviews as the next largest edition, English Wikipedia dominates global knowledge access relative to other language editions. Readers are prone to assuming English Wikipedia as a superset of all language editions, leading many to prefer it even when their primary language is not English. Other language editions, however, comprise complementary facts rooted in their respective cultures and media environments, which are marginalized in English Wikipedia. While Wikipedia’s user interface enables switching between language editions through its Interlanguage Link (ILL) system, it does not reveal to readers that other language editions contain valuable, complementary information. We present WikiGap, a system that surfaces complementary facts sourced from other Wikipedias within the English Wikipedia interface. Specifically, by combining a recent multilingual information-gap discovery method (Samir et al., 2024) with a user-centered design, WikiGap enables access to complementary information from French, Russian, and Chinese Wikipedia. In a mixed-methods study (n=21), WikiGap significantly improved fact-finding accuracy, reduced task time, and received a 32-point higher usability score relative to Wikipedia’s current ILL-based navigation system. Participants reported increased awareness of the availability of complementary information in non-English editions and reconsidered the completeness of English Wikipedia. WikiGap thus paves the way for promoting epistemic equity across language editions. Taken together, these findings show how interface design can redistribute reader attention and engagement across Wikipedia’s multilingual knowledge production communities, supporting epistemic equity.
@@ 1. Introduction
Wikipedia hosts more than 300 language editions, inviting contributions from many distinct linguistic and geographic regions. The investment in creating multiple language editions reflects Wikipedia’s principles of democratizing both access to and contribution of information. These language editions are sociogeographically situated; it is generally well understood in Wikipedia scholarship that they are not mere translations of one another (Hecht, 2013). Instead, the language editions tend towards reflecting the knowledge that is contextually relevant, often based on cultural experience and localized media diets (Avieson and Mazumdar, 2025; Kumar, 2021).
However, this heterogeneity is not well understood by the broader public that engage with Wikipedia, whether readers or contributors (Miquel-Ribé and Laniado, 2018). Rather, the popular understanding of multilingual Wikipedia is that there is a global consensus on what information is reliable and noteworthy (Hecht, 2013). Moreover, English Wikipedia is by and large perceived as the most comprehensive and complete account of this consensus, containing as much as or more information than any other, making other languages the minority. This English-as-superset assumption is widespread, even among people who speak multiple languages proficiently; in a 2011 report, $93\%$ of all editors surveyed attested to reading English Wikipedia, even though it was a primary language for only $52\%$ of them (Kumar, 2021; Wikimedia Foundation, 2011). In the same vein, $20\%$ of English Wikipedia’s visitors were based in India, even more than readers from the United Kingdom. English Wikipedia is so widely understood to be an epistemic authority that is commonly treated as a source of neutral and complete information in and of itself by editors contributing to minority language editions (Hickman et al., 2021).
This epistemic assumption is consequential not only for readers’ understanding of Wikipedia but also for the organization of its peer production communities. Wikipedia is sustained by interconnected yet unevenly resourced volunteer communities, each centered on different language editions (Sen et al., 2015). Prior research has shown that volunteer engagement in Wikipedia depends on motivational factors such as enjoyment, opportunities for learning, and perceived recognition within a contributor community, beyond shared ideological commitments to free and open knowledge (Nov, 2007; Arazy et al., 2017). When English Wikipedia is widely perceived as the most complete and authoritative source, reader attention and contributor effort become disproportionately concentrated in the English-language community. This concentration structurally amplifies the conditions that prior work has shown to support sustained participation, while making those conditions less accessible in minority language editions. One byproduct of this imbalance is a reduction in incentives for volunteer contributions to those communities (Maiberg, 2025), undermining Wikipedia’s ethos as a globally and cooperatively produced multilingual knowledge infrastructure. In this work, we ask how interface design can redistribute attention and engagement across Wikipedia’s multilingual peer production communities, promoting epistemic equity.  [^1] Throughout this paper, we use epistemic equity as a design lens, focusing specifically on how interface design can redistribute attention and engagement across Wikipedia’s multilingual peer production communities, rather than addressing broader institutional or political dimensions of epistemic equity.
Prior works have sought to incorporate contributions in minority language editions by consolidating information from all editions into a single unified knowledge base (e.g., Duh et al., 2013; Adar et al., 2009). However, this conceptual formulation fails to center minority language editions and the unique volunteer contributions they contain. It continues to perpetuate the ideology of there being a global consensus on noteworthy and reliable information, a universal knowledge base. As such, it does little to challenge the status quo of English Wikipedia as an objective and complete epistemic authority.
@@
In our work we take a social constructionist view of knowledge (Haraway, 2013) in approaching the problem by improving engagement with minority language editions. That is, we center necessarily incomplete and sociogeographically positioned curation of knowledge. To highlight the positionality (rather than objectivity) of English Wikipedia, we aim to highlight its content gaps relative to other language editions. At the same time, we balance this intervention against being minimally disruptive to the current Wikipedia reading design (Grudin, 1988). We present WikiGap, an augmentation of the Wikipedia reading experience that surfaces sentence-level factual differences from English to 3 other language editions of Wikipedia (French, Russian, and Chinese). The system combines a computational backend (InfoGap) that identifies multilingual content gaps (Samir et al., 2024) with a browser-based interface that presents these differences through in-page highlights and a sidebar display (Figure 1).
@@
We evaluated WikiGap through a mixed-methods user study using a fact-finding task focused on culturally rich food articles – a domain where cross-lingual disparities are particularly salient (Laufer et al., 2015). Our results show that participants strongly preferred WikiGap over Wikipedia’s default interlanguage links system (ILLs), with System Usability Scale (SUS) scores 32 points higher on average. In terms of performance, participants were significantly more accurate and faster at answering factual questions when using WikiGap, suggesting that surfacing multilingual information directly within the English interface not only improves user satisfaction but also enhances learning efficiency.
@@
Through a Post-Study interview, we find that participants, prior to using WikiGap, were largely unaware of content gaps in English Wikipedia, demonstrating that the problematic English-as-superset epistemic assumption persists more than a decade after it was identified in seminal studies by Hecht (2013). Participants were overwhelmingly appreciative of WikiGap’s capacity to surface these substantive information gaps during the reading of English Wikipedia pages, and expressed openness and enthusiasm about interacting with multilingual content in English Wikipedia. Moreover, a few participants with prior editing experience described how the system could support their editorial workflows by enabling source verification and facilitating the integration of multilingual content. These findings underscore WikiGap’s broader value as a bridge between communities of editors and readers across languages.
In Section 6, we reflect on how WikiGap challenges the English-as-superset assumption by redistributing reader attention and engagement across Wikipedia’s multilingual peer production communities. We further interpret the system as a boundary object (Star and Griesemer, 1989; Star, 2010) that supports articulation work between readers and editors across language editions. We contribute a validated system design for surfacing fact-level cross-lingual information gaps in Wikipedia. The system combines a computational pipeline that detects and integrates missing facts from non-English editions with an unobtrusive interface for English readers. A mixed-methods evaluation (n=21) shows that this design improves fact-finding accuracy and efficiency, outperforms Wikipedia’s interlanguage-link system in usability, and increases engagement with multilingual knowledge—challenging assumptions about English Wikipedia’s epistemic completeness. [^2] Code and data will be made publicly available upon publication.
@@ 2. Background and Related Work
Wikipedia’s multilingual architecture aspires to democratize knowledge access across languages, yet its design and community practices systematically privilege English content. In §2.1, we examine how this “English-as-superset” assumption (Hecht, 2013) persists despite empirical evidence to the contrary, and how current Wikipedia interface reinforces epistemic hierarchies that undermine the platform’s multilingual vision by concentrating attention and participation within a single language edition. In §2.2, we explain that cross-linguistic variation should be surfaced rather than obscured by drawing literature from prior CSCW research on asset-based design (Wong-Villacres et al., 2020; Kretzman and McKnight, 1993; Hardy and Thebault-Spieker, 2024) and feminist epistemology (Haraway, 2013; Bardzell, 2010). Then, we review prior technical interventions designed to surface cross-linguistic knowledge differences, evaluating them along two critical dimensions: the granularity of differences they expose and their integration with readers’ existing browsing practices to support engagement across language communities. This analysis reveals persistent limitations that motivate our WikiGap design.
2.1. English-as-Superset and Epistemic Inequity in Multilingual Wikipedia
Wikipedia embodies a central tension in collaborative knowledge systems: although it is architecturally multilingual and aspires to provide knowledge to “every single person on the planet in their own language” (Cohen, 2008), its design and community practices systematically privilege English content (Foucault, 2013; Said, 1977). This imbalance shapes how knowledge is organized and encountered on the platform, influencing both which perspectives become visible and how reader attention is distributed across language-specific peer production communities. At the interface level, this dynamic obscures cross-linguistic differences and reinforces what Hecht (2013) term the “English-as-superset” assumption—the belief that English Wikipedia exhaustively represents the knowledge available across other language editions.
CSCW scholarship on epistemic equity (Ajmani et al., 2024; Benjamin, 2023) helps explain why this assumption matters: collaborative systems should support the visibility and engagement of diverse epistemic standpoints rather than subsuming them under a dominant perspective. In Wikipedia’s multilingual context, this means recognizing that language communities produce knowledge shaped by distinct cultural, geographic, and historical conditions that are not uniformly visible across language editions
2.1.1. English Wikipedia Is Treated as the Default, Yet Knowledge Gaps Persist
Given English’s status as a global lingua franca, the English edition of Wikipedia has far outpaced other language editions in size and coverage. As a result, it is often implicitly treated as a superset of their knowledge. Hecht (2013) identified this “English-as-superset” viewpoint, which, though rarely stated explicitly, remains prevalent—even among researchers who rely on multilingual Wikipedia to train text-generation and text-embedding models. However, substantial knowledge gaps persist between English and other language editions. Hecht and Gergle (2010) showed that English Wikipedia lacks coverage on many topics documented in other languages, while more recent work demonstrates that even when topics overlap, English articles often omit facts present elsewhere (Samir et al., 2024).
2.1.2. From Knowledge Gaps to Epistemic Invisibility
From an epistemic equity perspective, such variation is not an error or deficiency, but an expected outcome of situated knowledge production. One interpretation of Wikipedia’s multilingual vision assumes a language-agnostic set of facts that can be uniformly translated across languages—a view that treats language as a neutral vessel for meaning, abstracted from socio-geographic context (Golumbia, 2009, Chapter 4). Yet much of Wikipedia’s content concerns culturally and geographically situated histories, practices, and identities (Wikimedia Foundation, 2024). Accordingly, language editions are shaped by the communities that produce them (Johnson and Lescak, 2022) and exhibit a well-documented “self-focus bias” toward regionally relevant topics (Hecht and Gergle, 2009; Samir et al., 2024; Callahan and Herring, 2011). This emphasis on positionality echoes feminist epistemological arguments that knowledge is always produced from particular social, cultural, and linguistic standpoints (Haraway, 2013). When such epistemic differences are rendered invisible, they not only obscure situated knowledge but also limit opportunities for readers and editors to encounter and engage with the communities that produce it.
The “English-as-superset” assumption can thus contribute to cycles of epistemic invisibility (Ajmani et al., 2024), in which visibility and legitimacy are recursively reinforced within collaborative systems. On Wikipedia, these dynamics manifest in both editorial and readership practices. Bilingual editors often consult English Wikipedia to determine what content belongs in other language editions, treating it as a reference point (Hickman et al., 2021). Wikipedia’s emphasis on verifiability further privileges English-language sources, creating structural barriers for knowledge rooted in other linguistic or cultural contexts (Sen et al., 2015). Although multilingual editors can bridge across editions, they constitute only a small fraction of Wikipedia’s editor base (Hale, 2014).
Readers similarly assume that different language editions contain equivalent information (Section 5). As engagement with smaller editions declines, editors in those communities receive less visibility and recognition, weakening key motivations for participation (Nov, 2007; Arazy et al., 2017). This feedback loop reduces incentives for contribution and undermines Wikipedia’s vision as a collaboratively produced multilingual knowledge infrastructure (Maiberg, 2025).
2.1.3. Interface Design and the Obscuring of Epistemic Difference
These epistemic hierarchies are not only reflected in community norms and editorial practices, but are also enacted through interface design. Wikipedia primarily supports cross-lingual navigation through Interlanguage Links (ILLs), presented as a drop-down menu on each article. While ILLs enable access to other language editions, the interface provides no cues indicating that articles may differ substantially in content. By framing language switching as a matter of translation rather than epistemic difference, the design implicitly suggests a shared underlying knowledge base, limiting opportunities for readers to engage with the work of other language communities. Yet prior work shows that many unique facts remain distributed across non-local language editions, even for regionally associated topics (Samir et al., 2024). As a result, Wikipedia’s current interface renders epistemic differences structurally invisible to readers, making it difficult to engage with other language editions.
Together, this body of work motivates a focus on design interventions that challenge the “English-as-superset” assumption by making cross-linguistic differences visible and supporting engagement across Wikipedia’s multilingual peer production communities.